### v1

@@ Abstract
We compute new  exact analytic expressions for one-loop scalar effective actions in Kerr (A)dS black hole (BH) backgrounds in four and five dimensions. These are computed by the connection coefficients of the Heun equation via a generalization of the Gelfand-Yaglom formalism to second-order linear ODEs with regular singularities. The expressions we find are in terms of Nekrasov-Shatashvili special functions, making explicit the analytic properties of the one-loop effective actions with respect to the gravitational parameters and the precise contributions of the quasi-normal modes. The latter arise via an associated integrable system. In particular, we prove asymptotic formulae for large angular momenta in terms of hypergeometric functions and give a precise mathematical meaning to Rindler-like region contributions. Moreover we identify the leading terms in the large distance expansion as the point particle approximation of the BH and their finite size corrections as encoding the BH tidal response. We also discuss exact properties of the thermal version of the BH effective actions. Although we focus on the real scalar field in dS-Kerr and (A)dS-Schwarzschild in four and five dimensions, similar formulae can be given for higher spin matter and radiation fields in more general gravitational backgrounds.
@@ 1 Introduction
It would be interesting to investigate higher spin perturbations within this method [^3] See Grewal:2022hlo for a discussion about one-loop effective actions in this case.. Perturbations up to spin two of the four-dimensional Kerr black hole were studied in Bonelli:2022ten . These results can be used to describe the one-loop Euclidean gravitational partition function and complement the analysis of logarithmic corrections to Kerr BH thermodynamics recently studied in Kapec:2023ruw . We remark that one-loop effective actions in supergravity were used to compute logarithmic corrections to the entropy of supersymmetric BHs Banerjee:2010qc ; Sen:2014aja . By complementing the analysis performed in this paper with the study of the zero-modes of the one-loop operators one should be able to compute logarithmics corrections to the entropy of Kerr-(A)dS BHs.
@@
Other gravitational backgrounds can be analyzed with the same methods, see Bianchi:2021mft ; Bianchi:2021xpr ; Bianchi:2021yqs ; Bianchi:2022qph ; Bianchi:2022wku ; Bianchi:2023lrg ; Bianchi:2023rlt ; Bianchi:2023sfs ; Bianchi2021MoreOT ; Amado:2021erf ; Imaizumi:2022qbi ; Giusto:2023awo for QNMs analysis, and it should therefore be possible to extend our method to the study of one-loop effective actions also in these cases (see also Zhao:2024rrw ). Related parallel investigations on QNMs appear in Bucciotti:2023ets ; Lei:2023mqx ; Grozdanov:2023tag ; Hatsuda:2023geo ; Hatsuda:2021gtn ; Kehagias:2022ndy . As the results in the realm of micro-state counting have been encouraging, these should also apply to the framework considered in Krishnan:2023jqn . On the other hand, effective near-horizon symmetries have been analyzed in Charalambous:2022rre ; Charalambous:2024tdj and it could be interesting to consider them in the light of our results.
@@ 2 One-loop black hole effective actions and Gelfand-Yaglom theorem
Let us denote with $\psi^{(\hat{z})}_{1,\lambda}(z)$ the solution selected bye bye  the boundary condition at $z=\hat{z}$. Using the connection formulae, we can write
@@ 5.1 Wick rotation and the thermal version
where the substitution (101) is implied, and the structure depending on QNMs can be written from the second line.
@@ C.3 Regularized determinant for Heun differential operators
Comparing the differential operator $\mathcal{D}_{1}$ in (174) with the operator $\tilde{\mathcal{D}}$ in (172), we have that the two are related by the dictionary
@@
$\mathrm{det}\mathcal{D}=$ $\sum_{\theta^{\prime}=\pm}\frac{2\pi\Gamma(-2\theta^{\prime}v)\Gamma(1-2\theta^{\prime}v)}{\prod_{\sigma=\pm}\Gamma\left(\frac{1}{2}+a_{0}-\theta^{\prime}v+\sigma\,a_{t}\right)\Gamma\left(\frac{1}{2}-\theta^{\prime}v+a_{1}+\sigma\,a_{\infty}\right)}t^{a_{0}+\theta^{\prime}v}\exp(\frac{1}{2}\partial_{a_{0}}F(t)+\frac{1}{2}\partial_{a_{1}}F(t)-\frac{\theta^{\prime}}{2}\partial_{v}F(t)).$ (187)
@@ Appendix E Gauge theory conventions
The gauge parameter $a$ is expressed in a series expansion in the instanton counting parameter $t$, obtained by inverting the Matone relation

### v2

@@ Abstract
We compute new exact analytic expressions for one-loop scalar effective actions in Kerr (A)dS black hole (BH) backgrounds in four and five dimensions. These are computed by the connection coefficients of the Heun equation via a generalization of the Gelfand-Yaglom formalism to second-order linear ODEs with regular singularities. The expressions we find are in terms of Nekrasov-Shatashvili special functions, making explicit the analytic properties of the one-loop effective actions with respect to the gravitational parameters and the precise contributions of the quasi-normal modes. The latter arise via an associated integrable system. In particular, we prove asymptotic formulae for large angular momenta in terms of hypergeometric functions and give a precise mathematical meaning to Rindler-like region contributions. Moreover, we identify the leading terms in the large distance expansion as the point particle approximation of the BH and their finite size corrections as encoding the BH tidal response. We also discuss the exact properties of the thermal version of the BH effective actions by providing a proof of the DHS formula and explicitly computing it for new relevant cases. Although we focus on the real scalar field in dS-Kerr and (A)dS-Schwarzschild in four and five dimensions, similar formulae can be given for higher spin matter and radiation fields in more general gravitational backgrounds.
@@ 1 Introduction
It would be interesting to investigate higher spin perturbations within this method [^3] See Grewal:2022hlo for a discussion about one-loop effective actions in this case.. Perturbations up to spin two of the four-dimensional Kerr black hole were studied in Bonelli:2022ten . These results can be used to describe the one-loop Euclidean gravitational partition function and complement the analysis of logarithmic corrections to Kerr and Kerr-AdS BH thermodynamics recently studied in Kapec:2023ruw ; Maulik:2024dwq . For recent studies on quantum effects and their consequence on the thermodynamic behavior of near-extremal black holes see also Kapec:2024zdj ; Kolanowski:2024zrq . We remark that one-loop effective actions in supergravity were used to compute logarithmic corrections to the entropy of supersymmetric BHs Banerjee:2010qc ; Sen:2014aja . By complementing the analysis performed in this paper with the study of the zero-modes of the one-loop operators one should be able to compute logarithmics corrections to the entropy of Kerr-(A)dS BHs.
@@
Other gravitational backgrounds can be analyzed with the same methods, see Bianchi:2021mft ; Bianchi:2021xpr ; Bianchi:2021yqs ; Bianchi:2022qph ; Bianchi:2022wku ; Bianchi:2023lrg ; Bianchi:2023rlt ; Bianchi:2023sfs ; Bianchi2021MoreOT ; Amado:2021erf ; Imaizumi:2022qbi ; Giusto:2023awo ; Cipriani:2024ygw for QNMs analysis, and it should therefore be possible to extend our method to the study of one-loop effective actions also in these cases (see also Zhao:2024rrw ). Related parallel investigations on QNMs appear in Bucciotti:2023ets ; Lei:2023mqx ; Grozdanov:2023tag ; Hatsuda:2023geo ; Hatsuda:2021gtn ; Kehagias:2022ndy . As the results in the realm of micro-state counting have been encouraging, these should also apply to the framework considered in Krishnan:2023jqn . On the other hand, effective near-horizon symmetries have been analyzed in Charalambous:2022rre ; Charalambous:2024tdj and it could be interesting to consider them in the light of our results.
@@ 2 One-loop black hole effective actions and Gelfand-Yaglom theorem
Let us denote with $\psi^{(\hat{z})}_{1,\lambda}(z)$ the solution selected by the boundary condition at $z=\hat{z}$. Using the connection formulae, we can write
@@ 5.1 Wick rotation and the thermal version
where the substitution (101) is implied, and the structure depending on QNMs can be read from the second line.
@@ C.3 Regularized determinant for Heun differential operators
Comparing the differential operator $\mathcal{D}_{1}$ in (174) with the operator $\tilde{\mathcal{D}}$ in (172), we have that the two are related by the dictionary [^10] In (185) the parameter $a$ denotes the parameter of the Hypergeometric equation, whereas in (187) it denotes the v.e.v. parameter.
@@
$\mathrm{det}\mathcal{D}=$ $\sum_{\theta^{\prime}=\pm}\frac{2\pi\Gamma(-2\theta^{\prime}a)\Gamma(1-2\theta^{\prime}a)}{\prod_{\sigma=\pm}\Gamma\left(\frac{1}{2}+a_{0}-\theta^{\prime}a+\sigma\,a_{t}\right)\Gamma\left(\frac{1}{2}-\theta^{\prime}a+a_{1}+\sigma\,a_{\infty}\right)}t^{a_{0}+\theta^{\prime}a}\exp(\frac{1}{2}\partial_{a_{0}}F(t)+\frac{1}{2}\partial_{a_{1}}F(t)-\frac{\theta^{\prime}}{2}\partial_{a}F(t)).$ (187)
@@ Appendix E Gauge theory conventions
The gauge parameter $a$ is expressed in a series expansion in the instanton counting parameter $t$, obtained by inverting the Matone relation Matone:1995rx